# Ramsey Numbers in Kneser Graphs

### v1

@@ 1 Introduction
It is easy to see that $\beta\leq 3$. Indeed, let $S_{1},\dots,S_{6}\in{[3k]\choose k}$ be six sets with the property that for every $\in[3k], tt$ belongs to exactly two of the sets $S_{1},\dots,S_{6}$. (It is easy to construct such sets $S_{1},\dots,S_{6}$: take the six distinct sets in two disjoint perfect matchings in the $k$-uniform hypergraph on vertex set $[3k]$.) Now consider a red/blue edge-coloring of the complete graph on vertices $S_{1},\dots,S_{6}$. Since $R(3,3)=6$, this graph contains a monochromatic triangle $S_{i}S_{j}S_{k}$, and $S_{i}\cap S_{j}\cap S_{k}=\emptyset$ for every distinct $i,j,k\in[6]$.
In this paper we make progress on Problems 2 and 4 in two different ways. First, in Section 2 we improve the upper bound on $\beta$ to $7/3$.
@@
For any $0<\alpha<\frac{1}{4}$, there exists some large enough $n$ such that any red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle.
Note that for $\alpha=\frac{1}{4}$, for $n=8$ there exists a red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ with no monochromatic triangle, as implied by Theorem 12.
@@
Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). We have
$R^{\textrm{KG}}_{r}(s,t)\leq(M+1)r+m-1.$
In particular,
@@
In Section 3 we establish improved upper bounds on $R^{\textrm{KG}}_{r}(s,t)$ for certain families of $s,t$.
@@
Using a computer code, we also obtain an exact value on $R^{\textrm{KG}}_{2}(3,3)$. The computational proof of the next theorem is discussed in Subsection 3.3.
@@
We have $R^{\textrm{KG}}_{2}(3,3)=9$ and $R^{\textrm{KG}}_{3}(3,3)\leq 13$.
Some of our results are summarized in Table 1 (for $r=2$) and Table 2 (for $r=3$). In each cell of the tables, the first number listed is the lower bound and the second is the upper bound. The subscripts on the numbers are the theorems in which we obtain the bound. For example, we have $11\leq R^{\textrm{KG}}_{2}(3,4)\leq 16$, where the lower bound follows from Theorem 10 and the upper bound is proven in Theorem 8. If there is only one number, then it is an exact value. Note that $R^{\textrm{KG}}_{r}(s,t)=R^{\textrm{KG}}_{r}(t,s)$.
@@
We conclude this paper with some progress on a Ramsey-type question that was posted by Dömötör Pálvölgyi on mathoverflow in 2016 [6]. Here an induced $\textrm{KG}(3r,r)$ is a subgraph of $\textrm{KG}(n,r)$ induced on vertex set $S\choose r$}, where $S\subset[n]$ is a set of size $3r$.
@@
Let $k_{1}\geq k_{2}\geq\dots\geq k_{t}\geq 2$ and suppose $n\geq k_{1}r+\sum_{i=2}^{t}(k_{i}-1)$. Let $F_{1},F_{2},\dots,F_{t}$ be a partition of $\binom{[n]}{r}$. Then, for some $i\in[t]$, $F_{i}$ contains $k_{i}$ pairwise disjoint sets.
Observe that when $k_{1}=k_{2}=\cdots=k_{t}=2$, Theorem 15 recovers the chromatic number of the Kneser graph. When $t=2$, we apply a simpler form.
Theorem 16 (Alon - Frankl - Lovász [1]).
Let $n\geq(r+1)k-1$ and let $F_{1},F_{2}$ be a partition of $\binom{[n]}{r}$. Then, either $F_{1}$ or $F_{2}$ contains $k$ pairwise disjoint sets.
@@ 2 The Holmsen - Hrusak - Roldán-Pensado problem
Our next goal is to prove Theorem 6, which gives a positive answer to Problem 2 in the case $\alpha<1/4$ and $t=3$. In order to prove the theorem, we need the following auxiliary result.
@@
Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). Let $0<\varepsilon<\frac{1}{M+1}$ and set $\alpha=\frac{1}{M+1}-\varepsilon$. Then, there exists some $n\in\mathbb{N}$ such that $\alpha n\in\mathbb{N}$ and $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$.
Proof of Theorem 17.
Fix $r\geq 2$ and set Nr=Mr+m1N-r=Mr+m-1. First, we will show that in any red/blue coloring of $G=\textrm{KG}(N,r)$, there is a red $K_{s}$ or a blue $K_{t}$. Suppose for contradiction that there is some coloring $c:E(G)\to\{R,B\}$ of $G$ without a red $K_{s}$ or blue $K_{t}$. Let $v=[N]\setminus[N-r]\in V(G)$. Then, $\binom{[N-r]}{r}=\binom{[Mr+m-1]}{r}$ can be partitioned into $N_{R}(v)$ and $N_{B}(v)$. Since $c$ contains no red $K_{s}$ and no blue $K_{t}$, then $G[N_{R}(v)]$ contains no red $K_{s-1}$ and no blue $K_{t}$. Similarly, $G[N_{B}(v)]$ contains no red $K_{s}$ and no blue $K_{t-1}$. In particular, $G[N_{R}(v)]$ contains no $K_{M}$ and $G[N_{B}(v)]$ contains no $K_{m}$. However, by Theorem 16, setting $F_{1}=N_{R}(v)$ and $F_{2}=N_{B}(v)$, either $N_{R}(v)$ contains a clique of size $M$ or $N_{B}(v)$ contains a clique of size $m$, a contradiction.
For the general case, fix $0<\varepsilon<\frac{1}{M+1}$ and set $\alpha=\frac{1}{M+1}-\varepsilon$. Choose $r\in\mathbb{Z}$ such that $\frac{r}{(M+1)r+m-1}\geq\alpha$. Set Nr=Mr+m1N-r=Mr+m-1 and $n=\frac{r}{\alpha}$. Then, by choice of $r$, observe that $N\leq n$. Therefore, since $\textrm{KG}(N,\alpha n)\subseteq\textrm{KG}(n,\alpha n)$, any red-blue coloring of $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$. ∎
Note that the general bound on $R^{\textrm{KG}}_{r}(s,t)$ in Theorem 7 is proven in the first paragraph of the proof of Theorem 17 above.
@@
We apply Theorem 17 with $s=t=3$. This gives $M=m=R(2,3)=3$. For any $0<\varepsilon<\frac{1}{4}$, we have $\alpha=\frac{1}{4}-\varepsilon$, and by Theorem 17 there exists some $n\in\mathbb{N}$ with $\alpha n\in\mathbb{N}$ such that any red/blue coloring of the edges of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle. ∎
@@ 3.1 Upper bounds
Let N=rR(s2,s)+2R(s2,s1)+2rN=rR(s-2,s)+2R(s-2,s-1)+2r, $G=\textrm{KG}(N,r)$ and let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$. For sake of contradiction, suppose there is no red $K_{s-1}$ nor blue $K_{s}$. If there is no blue edge in $G$, then we are done. So, set $u=\{N,N-1,\dots,N-r+1\}\in V(G)$ and $v=\{N-r,N-r-1,\dots,N-2r+1\}\in V(G)$. Without loss of generality, suppose $c(u,v)=B$. Then, $N(u)\cap N(v)\cong G^{\prime}=\textrm{KG}(N-2r,r)$. We will partition the vertices of $G^{\prime}$ into 4 sets. For $(X,Y)\in\{(R,R),(R,B),(B,R),(B,B)\}$, we denote by $N_{XY}$ the set of vertices $w$ in $G^{\prime}$ such that $c(w,u)=X$ and $c(w,v)=Y$.
Observe that $N_{BR}$ must contain no red $K_{s-2}$ or we may find a red $K_{s-1}$ along with $v$, and $N_{BR}$ must contain no blue $K_{s-1}$ or we may find a blue $K_{s}$ along with $u$. Thus $\omega(G^{\prime}[N_{BR}])\leq R(s-2,s-1)-1$. Similarly, $N_{RR}\cup N_{RB}$ must contain no red $K_{s-2}$ or blue $K_{s}$, so $\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq R(s-2,s)-1$. Furthermore, $N_{BB}$ must contain no red $K_{s-1}$ and no blue $K_{s-2}$, as otherwise we can find a blue $K_{s}$ along with the edge $uv$. Hence $\omega(G^{\prime}[N_{BB}])\leq R(s-2,s-1)-1$.
However, invoking Theorem 15 with $F_{1}=N_{RR}\cup N_{RB}$ ($k_{1}=R(s-2,s)$), $F_{2}=N_{BR}$ ($k_{2}=R(s-2,s-1)$), and $F_{3}=N_{BB}$ ($k_{3}=R(s-2,s-1)$), since $k_{1}\cdot r+(k_{2}-1)+(k_{3}-1)=N-2r$, either $F_{1}$ contains a clique on R(s2,s)R(s-2,s) vertices, $F_{2}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices, or $F_{3}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices. In any case, we are guaranteed either a red $K_{s-1}$ or a blue $K_{s}$, a contradiction. Therefore, $R^{\textrm{KG}}_{r}(s-1,s)\leq N$. ∎
Theorem 9 gives an improvement when $r=2$ and $s=3$.
@@
For sake of contradiction, fix N>2(R(3,t2))+2tN>2(R(3,t-2))+2t, let $G=\textrm{KG}(N,2)$, $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$, and suppose there is no red $K_{3}$ or blue $K_{t}$. Note that if there are no blue edges, then there must be a red $K_{3}$, and we have arrived at our contradiction.
Hence, without loss of generality, we may assume $c(\{1,2\},\{3,4\})=B$. Next, we look at $N(\{1,2\})\cap N(\{3,4\})\cong G^{\prime}=\textrm{KG}(N-4,2)$, and we will partition the vertices of $G^{\prime}$ into $4$ sets. For $(X,Y)\in\{(R,R),(R,B),(B,R),(B,B)\}$, we denote by $N_{XY}$ the set of vertices $v$ in $G^{\prime}$ such that $c(v,\{1,2\})=X$ and $c(v,\{3,4\})=Y$.
First, observe that for any two vertices $u,v\in N_{RB}$ with $uv\in E(G^{\prime})$, $c(u,v)=B$ (otherwise there is a red $K_{3}$ with $u,v,\{1,2\}$, a contradiction). Similarly, for each case of $u,v\in N_{RR}$, $u,v\in N_{BR}$, and $u\in N_{RR},v\in N_{RB}$, we have $c(u,v)=B$. These observations allow us to bound the clique numbers of $N_{RR}\cup N_{RB}$ and $N_{BR}$. In particular, $\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq t-1$ since all edges are blue and we cannot have a blue $K_{t}$. Additionally, $\omega(G^{\prime}[N_{BR}])\leq t-2$ since all edges are blue, and if there were a blue $K_{t-1}$, together with vertex $\{1,2\}$, we would have a blue $K_{t}$.
Finally, $\omega(G^{\prime}[N_{BB}])<R(t-2,3)$ since otherwise, there would be either a copy of a red $K_{3}$ (a contradiction), or a blue copy of $K_{t-2}$, which together with the blue edge $(\{1,2\},\{3,4\})$ would be a blue $K_{t}$, another contradiction.
By Theorem 15, with $F_{1}=N_{RR}\cup N_{RB}$ ($k_{1}=R(t-2,3))$, $F_{2}=N_{BR}$ ($k_{2}=t$), and $F_{3}=N_{BB}$ ($k_{3}=t-1$), since $N-4\geq 2(R(t-2,3))+(t-1)+(t-1-1)$, either $F_{1}$ contains a $K_{R(t-2,3)}$, $F_{2}$ contains a $K_{t}$, or $F_{3}$ contains a $K_{t-1}$, a contradiction in all cases. This final contradiction proves the result. ∎

### v2

@@ 1 Introduction
It is easy to see that $\beta\leq 3$. Indeed, let $S_{1},\dots,S_{6}\in{[3k]\choose k}$ be six sets with the property that for every $\in[3k], ii$ belongs to exactly two of the sets $S_{1},\dots,S_{6}$. (It is easy to construct such sets $S_{1},\dots,S_{6}$: take the six distinct sets in two disjoint perfect matchings in the $k$-uniform hypergraph on vertex set $[3k]$.) Now consider a red/blue edge-coloring of the complete graph on vertices $S_{1},\dots,S_{6}$. Since $R(3,3)=6$, this graph contains a monochromatic triangle $S_{i}S_{j}S_{k}$, and $S_{i}\cap S_{j}\cap S_{k}=\emptyset$ for every distinct $i,j,k\in[6]$.
In this paper we make progress on Problems 2-4 in two different ways. First, in Section 2 we improve the upper bound on $\beta$ in Problem 4 to $7/3$.
@@
For any rational number $\alpha$ such $0<\alpha<\frac{1}{4}$, there exists some large enough integer $n$ such that $\alpha n$ is an integer and any red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle.
Note that when $\alpha=\frac{1}{4}$, for $n=8$ there exists a red/blue edge-coloring of $\textrm{KG}(n,\alpha n)$ with no monochromatic triangle, as implied by Theorem 12 (an explicit coloring is given in Appendix A). We do not know if this holds for larger $n$.
@@
Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). Then
$R^{\textrm{KG}}_{r}(s,t)\leq(M+1)r+m-1,$
and in particular,
@@
In Section 3 we establish further upper bounds on $R^{\textrm{KG}}_{r}(s,t)$ for certain families of $s,t$.
@@
Using a computer code, we also obtain an exact value on $R^{\textrm{KG}}_{2}(3,3)$ and improve bounds on other Kneser Ramsey numbers. The computational proof of the next theorem is discussed in Subsection 3.3.
@@
We have
$R^{\textrm{KG}}_{2}(3,3)=9,~~R^{\textrm{KG}}_{3}(3,3)\leq 13,\text{~~and~~}12\leq R^{\textrm{KG}}_{2}(4,3)\leq 14.$
Some of our results are summarized in Table 1 (for $r=2$) and Table 2 (for $r=3$). In each cell of the tables, the first number listed is the lower bound and the second is the upper bound. The subscripts on the numbers are the theorems in which we obtain the bound. For example, we have $16\leq R^{\textrm{KG}}_{2}(3,5)\leq 23$, where the lower bound follows from Theorem 10 and the upper bound is proven in Theorem 9. If there is only one number, then it is an exact value. Note that $R^{\textrm{KG}}_{r}(s,t)=R^{\textrm{KG}}_{r}(t,s)$.
@@
We conclude this paper with some progress on a Ramsey-type question that was posted by Dömötör Pálvölgyi on mathoverflow in 2016 [6].
Given integer $n\geq s\geq r$, an induced $\textrm{KG}(s,r)$ is a subgraph $G$ of the graph $\textrm{KG}(n,r)$ induced on a vertex set $S\choose r$}, where $S\subset[n]$ is a set of size $s$. We say that $S$ is the ground set of $G$.
@@
Let $n\geq 2r\geq 2$ and $k_{1}\geq k_{2}\geq\dots\geq k_{t}\geq 2$ be integers, and suppose $n\geq k_{1}r+\sum_{i=2}^{t}(k_{i}-1)$. Let $F_{1},F_{2},\dots,F_{t}$ be a partition of $\binom{[n]}{r}$. Then, for some $i\in[t]$, $F_{i}$ contains $k_{i}$ pairwise disjoint sets.
Observe that when $k_{1}=k_{2}=\cdots=k_{t}=2$, Theorem 15 recovers the chromatic number of the Kneser graph.
@@ 2 The Holmsen - Hrusak - Roldán-Pensado problem
Our next goal is to prove Theorem 6, which gives a positive answer in Problem 2 in the case $\alpha<1/4$ and $t=3$. In order to prove the theorem, we need the following auxiliary result.
@@
Fix $s,t\geq 3$ and suppose $R(s-1,t)\geq R(s,t-1)$. Set M=R(s1,t)M=R(s-1,t) and m=R(s,t1)m=R(s,t-1). Let $\alpha$ be a rational number so that $0<\alpha<\frac{1}{M+1}$. Then, there exists some integer $n$ such that $\alpha n$ is an integer and $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$.
Proof.
Fix an integer $r\geq 2$ and let N=r(M+1)+m1N=r(M+1)+m-1. We first show that in any red/blue coloring of $G=\textrm{KG}(N,r)$, there is a red $K_{s}$ or a blue $K_{t}$. Indeed, assume for contradiction that there is some coloring $c:E(G)\to\{R,B\}$ of $G$ with no red $K_{s}$ or blue $K_{t}$. Let $v=[N]\setminus[N-r]\in V(G)$. Then, $\binom{[N-r]}{r}=\binom{[Mr+m-1]}{r}$ can be partitioned as $N_{R}(v)\cup N_{B}(v)$, where $N_{R}(v)$ and $N_{B}(v)$ are the red and blue neighborhoods of $v$, respectively. Since $c$ contains no red $K_{s}$ and no blue $K_{t}$, then $G[N_{R}(v)]$ contains no red $K_{s-1}$ and no blue $K_{t}$, and similarly, $G[N_{B}(v)]$ contains no red $K_{s}$ and no blue $K_{t-1}$. In particular, $G[N_{R}(v)]$ contains no $K_{M}$ and $G[N_{B}(v)]$ contains no $K_{m}$. However, by Theorem 15, setting $F_{1}=N_{R}(v)$ and $F_{2}=N_{B}(v)$, we have that either $N_{R}(v)$ contains a clique of size $M$ or $N_{B}(v)$ contains a clique of size $m$, a contradiction.
Now, let $\alpha$ be a rational number so that $0<\alpha<\frac{1}{M+1}$. Let $r$ be a positive integer so that $\frac{r}{(M+1)r+m-1}\geq\alpha$ and $\frac{r}{\alpha}$ is an integer. Let N=r+Mr+m1N=r+Mr+m-1 and $n=\frac{r}{\alpha}$. By the choice of $r$ we have $n\geq N$. Therefore, $\textrm{KG}(N,r)\subseteq\textrm{KG}(n,\alpha n)$, and thus any red-blue coloring of $\textrm{KG}(n,\alpha n)$ contains a red $K_{s}$ or a blue $K_{t}$. ∎
Note that the general bound on $R^{\textrm{KG}}_{r}(s,t)$ in Theorem 7 is proven in the first paragraph of the proof of Theorem 16.
@@
We apply Theorem 16 with $s=t=3$. This gives $M=m=R(2,3)=3$. Let $\alpha$ be a rational number satisfying $0<\alpha<\frac{1}{4}$. Then $\alpha<\frac{1}{M+1}$, and thus by Theorem 16 there exists some integer $n$ such that $\alpha n$ is an integer and any red/blue coloring of the edges of $\textrm{KG}(n,\alpha n)$ contains a monochromatic triangle. ∎
@@ 3.1 Upper bounds
The proofs of both Theorem 8 and Theorem 9 follow the same scheme, described as follows. For some integers $N,r,s,t$, let $G=\textrm{KG}(N,r)$ and let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$. Assume that $c$ has no red $K_{s}$ or blue $K_{t}$, and suppose there exists a blue edge in $c$. Then (by renaming the integers in $[N]$) we may assume without loss of generality that $c(uv)=B$, where $u=\{N,N-1,\dots,N-r+1\}\in V(G)$ and $v=\{N-r,N-r-1,\dots,N-2r+1\}\in V(G)$. Note that $G[N_{G}(u)]\cap G[N_{G}(v)]=\textrm{KG}(N-2r,r)$.
We now partition the set of vertices $N-2r]\choose r$} of $G^{\prime}=\textrm{KG}(N-2r,r)$ into four sets $N_{XY}$, where $X,Y\in\{R,B\}$, where $N_{XY}$ is the set of vertices $w\in{[N-2r]\choose r}$ such that $c(wu)=X$ and $c(wv)=Y$.
Observe that $N_{BR}$ contains no red $K_{s-1}$, for otherwise we have a red $K_{s}$ together with $v$, and additionally, $N_{BR}$ contains no blue $K_{t-1}$, for otherwise have a blue $K_{t}$ together with $u$. Thus
$\omega(G^{\prime}[N_{BR}])\leq R(s-1,t-1)-1$
where $\omega(H)$ is the clique number of the graph $H$. Similarly, $N_{RR}\cup N_{RB}$ contains no red $K_{s-1}$ or blue $K_{t}$, so
$\omega(G^{\prime}[N_{RR}\cup N_{RB}])\leq R(s-1,t)-1.$
Finally $N_{BB}$ contains no red $K_{s}$ and no blue $K_{t-2}$, as otherwise we can find a blue $K_{t}$ together with the edge $uv$. Hence
$\omega(G^{\prime}[N_{BB}])\leq R(s,t-2)-1.$
Now we would like to apply Theorem 15 to conclude that at least one of the above three induced subgraph of $G^{\prime}$ contains a large clique, which will constitute a contradiction. But in order to do that, we need to compare the sizes of the Ramsey numbers R(s1,t1)R(s-1,t-1), R(s1,t)R(s-1,t), and R(s,t2)R(s,t-2), which we do not know how to do in general. However, in the two special cases where $s$ and $t$ differ by 1 (Theorem 8) or when $s=3$ (Theorem 9) we can find the largest value among the three Ramsey numbers, which allows us to apply Theorem 15.
@@
Let N=rR(s2,s)+2(R(s2,s1)1)+2rN=rR(s-2,s)+2(R(s-2,s-1)-1)+2r, $G=\textrm{KG}(N,r)$ and let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$. Assume for contradiction that $c$ has no red $K_{s-1}$ or blue $K_{s}$. Since
$|V(G)|={N\choose r}\geq N>rR(s-2,s)=rR(s,s-2),$
there is a clique of size R(s,s2)R(s,s-2) in $G$, so if $c$ has no blue edge it must have a red $K_{s}$. So we may assume some edge is colored blue. Without loss of generality, suppose $c(uv)=B$, where $u=\{N,N-1,\dots,N-r+1\}\in V(G)$ and $v=\{N-r,N-r-1,\dots,N-2r+1\}\in V(G)$. Then $G[N_{G}(u)]\cap G[N_{G}(v)]=\textrm{KG}(N-2r,r)$. Let $G^{\prime}=\textrm{KG}(N-2r,r)$.
As explained above, we have (for $(s,t)\leftarrow(s-1,s)$)
$\begin{split}\omega(G^{\prime}[N_{BR}])&\leq R(s-2,s-1)-1,\\ \omega(G^{\prime}[N_{RR}\cup N_{RB}])&\leq R(s-2,s)-1,\\ \omega(G^{\prime}[N_{BB}])&\leq R(s-1,s-2)-1=R(s-2,s-1)-1.\end{split}$ (1)
Clearly, we have $R(s-2,s)\geq R(s-2,s-1)$, and thus we can invoke Theorem 15 with $F_{1}=N_{RR}\cup N_{RB}$, $F_{2}=N_{BR}$ and $F_{3}=N_{BB}$, and with $k_{1}=R(s-2,s),$ $k_{2}=R(s-2,s-1),$ $k_{3}=R(s-2,s-1)$. Since
$k_{1}\cdot r+(k_{2}-1)+(k_{3}-1)=N-2r,$
Theorem 15 entails that either $F_{1}$ contains a clique on R(s2,s)R(s-2,s) vertices, or $F_{2}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices, or $F_{3}$ contains a clique on R(s2,s1)R(s-2,s-1) vertices, contradicting (1). ∎
@@
We now apply our proof scheme with $(s,t)\leftarrow(3,t)$. We would like to show that for $t\geq 5$,
$R^{\textrm{KG}}_{r}(3,t)\leq rR(3,t-2)+2r+2t-3.$
Let N=rR(3,t2)+2r+2t3N=rR(3,t-2)+2r+2t-3 and $G=\textrm{KG}(N,r)$. Let $c:E(G)\to\{R,B\}$ be a red/blue edge-coloring of $G$. Assume for contradiction that $c$ has no red $K_{3}$ nor blue $K_{t}$. Like before, since N>rR(3,t2)N>rR(3,t-2), if $c$ has no blue edge, then it must have a red $K_{3}$. So we may assume some edge is colored blue, and without loss of generality, suppose $c(uv)=B$, where $u$ and $v$ as before. Then $G[N_{G}(u)]\cap G[N_{G}(v)]=\textrm{KG}(N-2r,r)$. Let $G^{\prime}=\textrm{KG}(N-2r,r)$.
Now we have
$\begin{split}\omega(G^{\prime}[N_{BR}])&\leq R(2,t-1)-1=t-2,\\ \omega(G^{\prime}[N_{RR}\cup N_{RB}])&\leq R(2,t)-1=t-1,\\ \omega(G^{\prime}[N_{BB}])&\leq R(3,t-2)-1.\end{split}$ (2)
It is straightforward to see that R(3,t2)>2(t3)R(3,t-2)>2(t-3) (indeed, take two disjoint cliques on t3t-3 vertices each, and color their edges blue; color the edges between them red; then the red graph is bipartite so contains no triangles). This implies that for $t\geq 5$ we have $R(3,t-2)\geq t$. Therefore, applying Theorem 15 with $F_{1}=N_{RR}\cup N_{RB}$, $k_{1}=R(3,t-2)$, $F_{2}=N_{BR}$, $k_{2}=t$, $F_{3}=N_{BB}$, and $k_{3}=t-1$, we have
$N-2r\geq r(R(3,t-2))+(t-1)+(t-1-1),$
implying that for some $i\in[3]$, $F_{i}$ contains a clique of size $k_{i}$, contradicting (2). ∎